Determine the missing amounts.

Unit Selling
Price

Unit Variable
Costs

Unit Contribution
Margin

Contribution Margin
Ratio

1. $600 $300 $enter a dollar amount (a) enter percentages % (b)
2. $300 $enter a dollar amount (c) $126 enter percentages % (d)
3. $enter a dollar amount (e) $enter a dollar amount (f) $690 30 %
Answer #1

Contribution Margin=Sales-Variable costs

Contribution Margin ratio=Contribution Margin/Sales

Unit selling price Unit variable cost Unit Contribution Margin Contribution Margin ratio
600 300 (600-300)=$300 (300/600)=50%
300 (300-126)=174 126 (126/300)=42%
(690/0.3)=$2300 (2300-690)=$1610 690 30%